OpenStudy (anonymous):

Find the exact value of the expression: cos^-1 (cos(-3π/5)) Explain please!

OpenStudy (anonymous):

cos and cos inverse cancel

OpenStudy (anonymous):

So the answer is -3π/5?

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

It said that the answer is 3π/5, without it being negative o:

OpenStudy (anonymous):

interesting

OpenStudy (anonymous):

How's that possible?

OpenStudy (chillout):

That's because cos(-x) = cos(x). You can check by doing cos(0-x) :)

OpenStudy (anonymous):

Oh, okay! Thank you guys

OpenStudy (chillout):

\[\cos (0-x) = \cos 0*\cos x + \sin 0*\sin x\]Thus,\[\cos (0 - x) = 1*\cos x + 0*\sin x\]So\[\cos(0 - x)=\cos x \leftarrow \rightarrow \cos(-x) = \cos(x)\]

OpenStudy (anonymous):

So does it work like that for sin, tan, cot, csc, and sec?

OpenStudy (chillout):

Not really. You gotta check the other ones by using trig identities. Here's a list: http://en.wikipedia.org/wiki/List_of_trigonometric_identities

OpenStudy (anonymous):

Alrighty, thanks!